# Michael Callen Miguel Fajardo-Steinhäuser Michael G. Findley

### v2

@@
Can Digital Aid Deliver During Humanitarian Crises? thanks: Callen: LSE, CERP and CESifo. Email: m.j.callen@lse.ac.uk. Fajardo-Steinhäuser: LSE. Email: m.fajardo-steinhauser@lse.ac.uk. Findley: University of Texas Austin. Email: mikefindley@utexas.edu. Ghani: Washington University in St. Louis. Email: tghani@wustl.edu. We thank Ilaha Eli Omar, Hila-Nawa Alam and the staff at Uplift Afghanistan Fund, Mukhtar Sabri and the staff at CDDO, Sanzar Kakar, Nigel Pont and the staff at HesabPay, and Kathleen McGowan for collaboration in program design and implementation. Cedric Antunes and Michael Cowan provided excellent research assistance. Shahim Kabuli provided excellent field management. Noam Angrist, Tim Besley, Robin Burgess, Michael Denly, Arif Husain, Daniel Nielson, Rohini Pande, Lamar Pierce, Soledad Prillaman, Simone Schaner, Jacob Shapiro, Abu Siddique, Garima Siwach, Rebecca Wolfe, Noam Yuchtman and seminar participants at the American Comparative Economic Society (ACES) political economy summer school, the Asian Development Bank Institute (ADBI) and Yale conference, Brookings, CEPR/TIME/TCD at Trinity College, HBS Tech for All Lab, Innovations for Poverty Action (IPA) / Northwestern, the Institute for Fiscal Studies (IFS), King’s College, John’s Hopkins School of Advanced International Studies (SAIS), The University of Maryland, Washington University in St. Louis, Wharton Management in Emerging Markets Conference, and the World Bank shared insightful feedback. Research funding was provided by the UK Foreign, Commonwealth & Development Office, awarded through the J-PAL Crime and Violence Initiative. The authors declare no competing interests. The research protocol was approved by the London School of Economics’ Institutional Review Board (study number 89546).
@@
(June 10, 2024)
Can digital payments systems help reduce extreme hunger? Humanitarian needs are at their highest since 1945, aid budgets are falling behind, and hunger is concentrating in fragile states where repression and aid diversion present major obstacles. In such contexts, partnering with governments is often neither feasible nor desirable, making private digital platforms a potentially useful means of delivering assistance. We experimentally evaluated digital payments to extremely poor, female-headed households in Afghanistan, as part of a partnership between community, nonprofit, and private organizations. The payments led to substantial improvements in food security and mental well-being. Despite beneficiaries’ limited tech literacy, 99.75% used the payments, and stringent checks revealed no evidence of diversion. Before seeing our results, policymakers and experts are uncertain and skeptical about digital aid, consistent with the lack of prior evidence on digital payments for humanitarian response. Delivery costs are under 7 cents per dollar, which is 10 cents per dollar less than the World Food Programme’s global figure for cash-based transfers. These savings can help reduce hunger without additional resources, demonstrating how hybrid partnerships utilizing digital platform technologies can help address grand challenges in difficult contexts.
Keywords: Hunger, Fragility, Digital payments, Grand challenges, Partnerships
@@ 1 Introduction
Can digital platform technologies enable more effective humanitarian crisis response? Humanitarians increasingly transfer physical cash to vulnerable individuals, but that modality is expensive and logistically challenging, subject to diversion by hostile actors that donors do not want to support, and requires in-person contact that is increasingly denied to marginalized populations. [^3] For example, the share of the World Food Program’s (WFP) assistance delivered as cash-based transfers (rather than in-kind aid contributions) rose from 2% in 2010 to 34% in 2021 (McDonough, 2022), a move supported by research evidence (World Bank, 2022). Relative to cash, digital payment systems create value by leveraging scalable technology and integrating private firms to facilitate transactions (cf. Dodgson et al., 2015; Wormald et al., 2021; Suri et al., 2023). Delivering aid digitally offers the potential to reduce coordination costs and delays, increase transparency for donors and preserve privacy of beneficiaries, and utilize local supply chains without relying on local authorities (Byrd, 2023; De Waal, 2023). While digital payment platforms are widely utilized by development programs for poverty alleviation (Muralidharan, Niehaus, and Sukhtankar, 2016; Suri and Jack, 2016), adoption by humanitarians has been slowed by market and institutional frictions in fragile states (Williamson, 2000; Dorobantu, Kaul, and Zelner, 2017; World Food Programme, 2021b) – particularly given limited evidence on the efficacy of digital aid to assist vulnerable, hard-to-reach populations (Aker et al., 2016; Pazarbasioglu et al., 2020; Gentilini, 2022). [^4] GiveDirectly (2024) catalogues the vast literature on cash transfers, including important contributions by Haushofer and Shapiro (2016); Bedoya Arguelles et al. (2019); Banerjee et al. (2020); Muralidharan et al. (2021); McIntosh and Zeitlin (2022). While prior cash transfer studies focus mainly on sustained poverty reduction (i.e. SDG 1), we are concerned with basic human survival (i.e. SDG 2); Online Appendix A discusses prior cash transfer studies focused on humanitarian goals.
@@
In this study, we conduct a randomized evaluation to study the impacts of a digital aid program in Afghanistan. Specifically, we compare a treatment group receiving digital aid payments against a comparison group which does not receive aid until later. The program was a hybrid partnership involving community, nonprofit, and private organizations. Locally elected community councils identified extremely poor, tech-illiterate, female-headed households in three cities: Kabul, Herat, and Mazar-i-Sharif. Working with an Afghan digital payments platform and a nationwide Afghan NGO, a U.S.-based nonprofit organization transferred digital value vouchers to beneficiaries’ mobile phones that could be spent at participating local private merchants. [^5] These value vouchers have a denominated local currency value redeemable at participating merchants for any available goods. World Food Programme (2023e) treats value vouchers as an official form of a cash-based transfer, unlike commodity vouchers, which are both tied to specific merchants and goods. The intervention transferred 45 USD payments every two weeks (the equivalent of 3.75 months worth of average monthly household income in the sample) for two months. The experiment included 2,409 beneficiary households, randomly assigned to treatment and control groups. Given our study’s focus on relieving human suffering, we make use of appropriate methods to measure key outcomes and to avoid false positives, including rigid adherence to a Pre-Analysis Plan (PAP). [^6] The Pre-Analysis Plan is registered here.
@@
Our work builds on two extensive bodies of social science research on cash transfers – one documenting the usefulness of digital transfers to achieve development outcomes (Bastagli et al., 2019; GiveDirectly, 2024) and the second focused mainly on the benefits of distributing physical cash to vulnerable populations during humanitarian crises (World Bank, 2022). While the former literature in development contexts often takes the existence of robust digital platforms for granted, the latter literature in humanitarian crises rarely engages the private sector as a delivery partner. In Online Appendix A we systematically review the intersection of these two literatures and find limited evidence regarding digital payments to hard-to-reach populations in fragile states during humanitarian crises – the closest digital aid papers examine comparatively stable settings including refugee camps and natural disaster response (cf. de Hoop, Morey, and Seidenfeld, 2019; MercyCorps, 2022). Our study demonstrates that digital platforms can establish trusted and secure payment channels linked to local retailers in fragile states, and that community and nonprofit partners can help facilitate adoption by the most vulnerable populations. In exploring the efficacy of digitization for humanitarian response, we echo the emphasis of George, Merrill, and Schillebeeckx (2021) on both technological innovation and “developing business models that infuse innovations with new purpose.” Indeed, the digital payments platform utilized in this study was developed prior to the Taliban’s 2021 return to power with the original intent of serving a growing Afghan middle class, and only afterwards shifted its focus to assist in aid delivery to the most vulnerable citizens.
@@ 2.1 Partners
The program was a hybrid arrangement designed to minimize interference by Taliban authorities and address common issues facing digital transfers – including technical issues like interoperability and smartphone ownership and social issues like tech literacy and trust. The research team composed of academics designed all research-related components of the project. Uplift Afghanistan, a U.S.-based nonprofit, initiated transfers to beneficiaries, which were identified and onboarded by an organisation of community-based institutions. Lastly, HesabPay, a digital payments platform that was compatible across mobile network operators, facilitated transactions through its network of private merchants. This division of labor is consistent with Luo and Kaul (2019), who argue for-profit firms are best-positioned to innovate (as in the case of building a digital payments ecosystem), self-governing collectives have a comparative advantage in private ordering (as in determining which potential beneficiaries to prioritize), and nonprofit organizations are well-suited for fiduciary roles.
@@
The private merchants enrolled in the digital payments platform’s acceptance network were pre-existing Afghan businesses providing basic foodstuffs and other household items for sale in the same neighborhoods as beneficiaries. Aid payments took the form of a digital value voucher denominated in local currency that could be exchanged for any available goods from any enrolled merchant. All beneficiaries conducted a test purchase with a merchant immediately following their onboarding session. While many chose to return to that specific merchant for future purchases, we confirm broader engagement with the acceptance network using beneficiaries’ transaction data from the digital payments provider. Table A1 reports the total number of merchants serving beneficiaries in the transaction data, which is consistently larger than the number from onboarding sessions. [^7] Note that this likely underestimates the true number of merchants in any given area, since i) we only contacted numbers with which our beneficiaries transacted more than 10 times, ii) among those numbers with which our participants transacted more than 10 times, we couldn’t reach all of them, and iii) there are more merchants in the areas that accept the digital vouchers but our participants did not visit. Figure A1 maps the locations of these merchants for each of the three cities in which we conducted the intervention, visually demonstrating the decentralized distribution network.
@@ 2.2 Beneficiary identification and onboarding
Locally-elected Community Development Councils (CDCs) identified female-headed households as program beneficiaries through community-level, participatory meetings. [^8] The CDCs were established through local elections as part of the National Solidarity Program starting in 2004 (Beath et al., 2016; Beath, Christia, and Enikolopov, 2017b; Beath, Christia, and Enikolopov, 2013, 2017a, 2018), where their primary job was to oversee block grants of development funding, and they were given a much broader range of local administrative authorities under the Citizens’ Charter, starting in 2016. Our goal was to identify about 2400 vulnerable women in three Afghan cities (Kabul, Herat and Mazar-i-Sharif) to be part of the intervention. To do so, we worked with the Community Driven Development Organization (CDDO), an Afghan organization that assists CDCs in a wide-array of local activities. Local CDCs identified potential beneficiaries through a well-being analysis in which community members, elders and mullahs together categorize all community households into different socioeconomic groups (e.g., well-off, middle income, poor, very poor). Our participants come from the lowest group. Thus, participants are identified through a process relying on the community deliberation regarding who is most vulnerable. We note that ownership of a feature phone was not a binding constraint for any beneficiaries, reflecting high nationwide levels of mobile phone ownership. [^9] Mobile phone ownership in Afghanistan has grown rapidly over the past two decades, from approximately 25,000 subscribers in 2002 to over 22 million subscribers in 2021 (World Bank, 2023b). In a nationally-representative survey, 91% of respondents reported at least one member of their household owned a mobile phone (66% of respondents report personally using a mobile phone), while 46% of that subgroup reported having an internet connection (Asia Foundation, 2019).
@@
After identification, participants were invited to local onboarding sessions where four activities took place. [^10] These took place during September 2022 in each of the three cities, with each session having between 24 and 80 participants. First, the program was described to them. Second, if they wanted to participate, their informed consent was collected and a baseline survey was administered. Third, each woman opened an account with HesabPay, an Afghan commercial digital payments platform. [^11] Launched in 2019, HesabPay is an interoperable digital payments platform that is feature-phone compatible and operates on the Algorand blockchain. At the time of the study, HesabPay had an active and growing local merchant network in Afghanistan’s major cities. Finally, each woman completed a test purchase with a nearby private merchant using the platform. [^12] During pilots we conducted prior to the actual intervention, it became clear that participants needed help at first using the digital payments platform, as most of them were illiterate and had no prior experience with mobile money. This is why we decided to use the onboarding sessions to both help participants create their accounts and resolve any remaining questions and also to have participants complete a test purchase with a nearby merchant using the mobile payment platform. The CDDO, working together with local CDCs, ensured the safety of staff and beneficiaries by informing local authorities and maintaining independence from external interference in the beneficiary selection process.

### v4

@@
Can Digital Aid Deliver During Humanitarian Crises? thanks: Callen: LSE, CERP and CESifo. Email: m.j.callen@lse.ac.uk. Fajardo-Steinhäuser: LSE. Email: m.fajardo-steinhauser@lse.ac.uk. Findley: University of Texas Austin. Email: mikefindley@utexas.edu. Ghani: Washington University in St. Louis. Email: tghani@wustl.edu. We thank Ilaha Eli Omar, Hila-Nawa Alam and the staff at Uplift Afghanistan Fund, Mukhtar Sabri and the staff at CDDO, Sanzar Kakar, Nigel Pont and the staff at HesabPay, and Kathleen McGowan for collaboration in program design and implementation. Cedric Antunes and Michael Cowan provided excellent research assistance. Shahim Kabuli provided excellent field management. Noam Angrist, Tim Besley, Robin Burgess, Michael Denly, Arif Husain, Daniel Nielson, Rohini Pande, Lamar Pierce, Soledad Prillaman, Simone Schaner, Jacob Shapiro, Abu Siddique, Garima Siwach, Rebecca Wolfe, Noam Yuchtman and seminar participants at the ACES summer school, the ADB and Yale conference, Brookings, CEPR/TIME/TCD at Trinity College, HBS Tech for All Lab, Innovations for Poverty Action/Northwestern, the Institute for Fiscal Studies, King’s College, John’s Hopkins School of Advanced International Studies, The University of Maryland, Washington University in St. Louis, Wharton Management in Emerging Markets Conference, and the World Bank shared insightful feedback. Research funding was provided by the UK Foreign, Commonwealth & Development Office, awarded through the J-PAL Crime and Violence Initiative. The authors declare no competing interests. The research protocol was approved by the London School of Economics’ Institutional Review Board (study number 89546).
@@
(October 28, 2024)
Can digital payments help reduce extreme hunger? Humanitarian needs are at their highest since 1945, aid budgets are falling behind, and hunger is concentrating in fragile states where repression and aid diversion present major obstacles. In such contexts, partnering directly with governments is often neither feasible nor desirable, making private digital payment platforms a potentially useful means of delivering assistance. We experimentally evaluated digital payments to extremely poor, female-headed households in Afghanistan, as part of a partnership between community, nonprofit, and private organizations. The payments led to substantial improvements in food security and mental well-being. Despite beneficiaries’ limited tech literacy, 99.75% used the payments, and stringent checks revealed no evidence of diversion. Before seeing our results, policymakers and experts are uncertain and skeptical about digital aid, consistent with the lack of prior evidence on digital payments for humanitarian response. Delivery costs are under 7 cents per dollar, which is 10 cents per dollar less than the World Food Programme’s global figure for cash-based transfers. These savings can help reduce hunger without additional resources, demonstrating how hybrid partnerships utilizing digital payment platforms can help address grand challenges in difficult contexts.
@@ 1 Introduction
Can digital payment platforms enable more effective humanitarian crisis response? Humanitarians increasingly transfer physical cash to vulnerable individuals, but that modality is expensive and logistically challenging, subject to diversion by hostile actors that donors do not want to support, and requires in-person contact that is increasingly denied to marginalized populations. [^3] For example, the share of the World Food Program’s (WFP) assistance delivered as cash-based transfers (rather than in-kind aid contributions) rose from 2% in 2010 to 34% in 2021 (McDonough, 2022), a move supported by research evidence (World Bank, 2022). Relative to cash, digital payment systems create value by leveraging scalable technology and integrating private firms to facilitate transactions (cf. Dodgson et al., 2015; Wormald et al., 2021; Suri et al., 2023). [^4] The United Nations Better than Cash Alliance defines a digital payment as ‘the transfer of value from one payment account to another using a digital device or channel. This definition may include payments made with bank transfers, mobile money, QR codes, and payment instruments such as credit, debit, and prepaid cards.’ In this study, we examine a specific digital payments platform in Afghanistan, HesabPay, that is available as an app on iOS and Android, and is also accessible to feature phone users via USSD or combined with a QR code card. Delivering aid digitally offers the potential to reduce coordination costs and delays, increase transparency for donors and preserve privacy of beneficiaries, and utilize local supply chains without relying directly on local authorities (Byrd, 2023; De Waal, 2023). While digital payment platforms are widely utilized by development programs for poverty alleviation (Muralidharan, Niehaus, and Sukhtankar, 2016; Suri and Jack, 2016), adoption by humanitarians has been slowed by market and institutional frictions in fragile states (Williamson, 2000; Dorobantu, Kaul, and Zelner, 2017; World Food Programme, 2021b) – particularly given limited evidence on the efficacy of digital aid to assist vulnerable, hard-to-reach populations (Aker et al., 2016; Pazarbasioglu et al., 2020; Gentilini, 2022). [^5] GiveDirectly (2024) catalogs the vast literature on cash transfers, including important contributions by Haushofer and Shapiro (2016); Bedoya Arguelles et al. (2019); Banerjee et al. (2020); Muralidharan et al. (2021); McIntosh and Zeitlin (2022). While prior cash transfer studies focus mainly on sustained poverty reduction (i.e. SDG 1), we are concerned with basic human survival (i.e. SDG 2); online Appendix A discusses prior cash transfer studies focused on humanitarian goals.
@@
In this study, we conduct a randomized evaluation to study the impacts of a digital aid program in Afghanistan. Specifically, we compare a treatment group receiving digital aid payments against a comparison group which does not receive aid until later. The program was a hybrid partnership involving community, nonprofit, and private organizations. Locally elected community councils identified extremely poor, tech-illiterate, female-headed households in three cities: Kabul, Herat, and Mazar-i-Sharif. Working with an Afghan digital payments platform and a nationwide Afghan NGO, a U.S.-based nonprofit organization transferred digital value vouchers to beneficiaries’ mobile phones that could be spent at participating local private merchants. [^6] These value vouchers have a denominated local currency value redeemable at participating merchants for any available goods. World Food Programme (2023e) treats value vouchers as an official form of a cash-based transfer, unlike commodity vouchers, which are both tied to specific merchants and goods. The intervention transferred 45 USD payments every two weeks (the equivalent of 3.75 months worth of average monthly household income in the sample) for two months. [^7] The Afghanistan Cash and Voucher Working Group (CVWG) estimates a Basic Food Basket for a family of seven costs approximately 96 USD per month (Bete, 2022), roughly equivalent to two biweekly direct aid transfers of 4000 AFA and a monthly survey incentive of 350 AFA. This basket was composed of 89 kg wheat flour, 21 kg domestic rice, 7 liters vegetable oil, 9 kg pulses, and 1 kg salt at prevailing exchange rates in August 2022. The experiment included 2,409 beneficiary households, randomly assigned to treatment and control groups. Given our study’s focus on relieving human suffering, we make use of appropriate methods to measure key outcomes and to avoid false positives, including rigid adherence to a Pre-Analysis Plan (PAP). [^8] The Pre-Analysis Plan is registered here. Crucially, the program leveraged three enabling conditions: widespread access to mobile phones, the digital platform’s robust merchant network, and sufficient market availability of food products to meet demand.
@@
Our work builds on two extensive bodies of social science research on cash transfers – one documenting the usefulness of digital transfers to achieve development outcomes (Bastagli et al., 2019; GiveDirectly, 2024) and the second focused mainly on the benefits of distributing physical cash to vulnerable populations during humanitarian crises (World Bank, 2022). While the former literature in development contexts often takes the existence of robust digital payment platforms for granted, the latter literature in humanitarian crises rarely engages the private sector as a delivery partner. In Online Appendix A, we systematically review the intersection of these two literatures and find limited evidence regarding digital payments to hard-to-reach populations in fragile states during humanitarian crises – the closest digital aid papers examine comparatively stable settings including refugee camps and natural disaster response (cf. de Hoop, Morey, and Seidenfeld, 2019; MercyCorps, 2022). Our primary contribution is to provide clear proof-of-concept that digital transfers can cost-effectively address humanitarian needs for highly vulnerable, hard-to-reach groups while avoiding diversion. In exploring the efficacy of digitization for humanitarian response, we echo the emphasis of George, Merrill, and Schillebeeckx (2021) on both technological innovation and “developing business models that infuse innovations with new purpose.” Indeed, the digital payments platform utilized in this study was developed prior to the Taliban’s 2021 return to power with the original intent of serving a growing Afghan middle class, and only afterwards shifted its focus to assist in aid delivery to the most vulnerable citizens.
@@ 2 Research design
2.1 Theory of change
In principle, digital payments should enable the transfer of value to vulnerable beneficiaries who can then purchase food and thus alleviate their household’s food insecurity. However, several enabling conditions must first be met: vulnerable households must have access to mobile phones, enough merchants must accept digital payments to allow convenient and competitive shopping, and the market must have enough goods to meet demand. Even if these conditions are met, other obstacles might still limit the effectiveness of digital aid: illiterate Afghan women may have difficulty utilizing digital payments, local merchants may find digital transactions overly cumbersome, and the Taliban authorities may decide to interfere.
Following an extensive piloting process, this digital aid program was designed with the goal of ensuring a successful implementation despite the challenges posed by the context and population. Beneficiaries were selected through locally-elected Community Development Councils (CDCs), who informed local Taliban authorities of their activities but denied them influence over participant selection. Tech illiteracy was addressed through a careful onboarding process described below, which also ensured that merchants were well-prepared to meet the needs of this population. Finally, research procedures were put in place to ensure high quality data and valid inference procedures to minimize the risk of false positives.
@@ 2.2 Partners
The program was a hybrid arrangement designed to minimize interference by Taliban authorities and address common issues facing digital transfers – including technical issues like interoperability and smartphone ownership and social issues like tech literacy and trust. The research team composed of academics designed all research-related components of the project. Locally-elected Community Development Councils (CDCs) identified female-headed households as program beneficiaries through community-level, participatory meetings and facilitated their engagement in the program. Uplift Afghanistan, a U.S.-based nonprofit, received grant funding from a private foundation and managed transfers to beneficiaries. Lastly, HesabPay, a digital payments platform that was compatible across mobile network operators, facilitated transactions through its network of private merchants. This division of labor is consistent with Luo and Kaul (2019), who argue for-profit firms are best-positioned to innovate (as in the case of building a digital payments ecosystem), self-governing collectives have a comparative advantage in private ordering (as in determining which potential beneficiaries to prioritize), and nonprofit organizations are well-suited for fiduciary roles. All partners shared a common interest in testing whether digital payments were a viable and attractive channel for humanitarian aid.
2.3 HesabPay’s technology and merchant network
HesabPay is a digital payments platform that is interoperable across Afghanistan’s mobile networks and transfers value using either a smartphone app or via a feature phone using a combination of transaction initiation via QR code and transaction verification via feature phone USSD. [^9] HesabPay was founded in 2016 by entrepreneur Sanzar Kakar. For more details, see the Algorand Foundation case study here. HesabPay uses the Algorand blockchain as a settlement layer, facilitating finalization and recording of digital transactions; HesabPay users are automatically registered for a custodial Algorand wallet (e.g. HesabPay holds their private keys on their behalf) and all transactions are automatically recorded on the Algorand blockchain without disclosing user identities. HesabPay is licensed by Afghanistan’s Central Bank as a financial service provider, similar to the regulatory framework for HesabPay’s main competition from Afghanistan’s mobile money operators.
@@ 2.3 HesabPay’s technology and merchant network
At the time of the study, HesabPay had an active and growing local merchant acceptance network in Afghanistan’s major cities, composed of over one thousand pre-existing Afghan businesses providing basic foodstuffs and other household items for sale in the same neighborhoods as beneficiaries. In this program, aid payments took the form of a digital value voucher denominated in local currency that could be exchanged for any available goods from any HesabPay enrolled merchant. All beneficiaries conducted a test purchase with a nearby merchant immediately following their onboarding session. While many chose to return to that specific merchant for future purchases, we confirm broader engagement with the acceptance network using beneficiaries’ transaction data from the digital payments provider. A1 reports the total number of merchants serving beneficiaries in the transaction data, which is consistently larger than the number from onboarding sessions. [^10] Note that this likely underestimates the true number of merchants in any given area, since i) we only contacted numbers with which our beneficiaries transacted more than 10 times, ii) among those numbers with which our participants transacted more than 10 times, we couldn’t reach all of them, and iii) there are more merchants in the areas that accept the digital vouchers but our participants did not visit. A1 maps the locations of these merchants for each of the three cities in which we conducted the intervention, visually demonstrating the decentralized distribution network.
@@ 2.4 Beneficiary identification and onboarding
Our goal was to identify about 2400 vulnerable women in three Afghan cities (Kabul, Herat and Mazar-i-Sharif) to be part of the intervention. To do so, we worked with the Community Driven Development Organization (CDDO), an Afghan organization that assists CDCs in a wide-array of local activities. [^11] The CDCs were established through local elections as part of the National Solidarity Program starting in 2004 (Beath et al., 2016; Beath, Christia, and Enikolopov, 2017b; Beath, Christia, and Enikolopov, 2013, 2017a, 2018), where their primary job was to oversee block grants of development funding, and they were given a much broader range of local administrative authorities under the Citizens’ Charter, starting in 2016. Local CDCs identified potential beneficiaries through a well-being analysis in which community members, elders and mullahs together categorize all community households into different socioeconomic groups (e.g., well-off, middle income, poor, very poor). Our participants come from the lowest group. Thus, participants are identified through a process relying on the community deliberation regarding who is most vulnerable. We note that ownership of a feature phone was not a binding constraint for any beneficiaries, reflecting high nationwide levels of mobile phone ownership. [^12] Mobile phone ownership in Afghanistan has grown rapidly over the past two decades, from approximately 25,000 subscribers in 2002 to over 22 million subscribers in 2021 (World Bank, 2023b). In a nationally-representative survey, 91% of respondents reported at least one member of their household owned a mobile phone (66% of respondents report personally using a mobile phone), while 46% of that subgroup reported having an internet connection (Asia Foundation, 2019).
@@
After identification, participants were invited to local onboarding sessions where four activities took place. [^13] These took place during September 2022 in each of the three cities, with each session having between 24 and 80 participants. First, the program was described to them. Second, if they wanted to participate, their informed consent was collected and a baseline survey was administered. Third, each woman opened an account with HesabPay, an Afghan commercial digital payments platform. Finally, each woman completed a test purchase with a nearby private merchant using the platform. [^14] During pilots we conducted prior to the actual intervention, it became clear that participants needed help at first using the digital payments platform, as most of them were illiterate and had no prior experience with mobile money. This is why we decided to use the onboarding sessions to both help participants create their accounts and resolve any remaining questions and also to have participants complete a test purchase with a nearby merchant using the mobile payment platform. The CDDO, working together with local CDCs, ensured the safety of staff and beneficiaries by informing local authorities and maintaining independence from external interference in the beneficiary selection process.